OpenStudy (anonymous):

G and H are mutually exclusive events. P(A)= 0.5; P(B)=0.3 Explain why this statement must be false: P(B I A)=0.4

OpenStudy (agent0smith):

If two events are mutually exclusive, then the probability of one event does NOT depend on the other occurring (or not occurring). ie P(B I A)=P(B) and P(A | B)=P(A)

OpenStudy (anonymous):

why must the following statement be false then?

OpenStudy (agent0smith):

Have a look at the probability values they gave you. P(B I A)=0.4 and P(B)=0.3 Now compare them to this... P(B I A)=P(B) which is true for mutually exclusive events.

OpenStudy (anonymous):

Ah, I see that now! Thank you.

OpenStudy (agent0smith):

You're welcome :)